A canola plant with genotype Bb is selfed and its progeny are distributed as as follows.

Genotype Number of individuals
BB 100
Bb 75
bb 10
Total 185

For A-E, round answers to three decimal places, and include a leading zero (ex. 0.150 or 0.151).

A. What is the frequency of the BB genotype?

B. What is the frequency of the Bb genotype?

C. What is the frequency of the bb genotype?

D. What is the frequency of the B allele?

E. What is the frequency of the b allele?
Answer #1

Answer ) To calculate the frequency of the genotypes, you needi to divide the individuals with that genotype divided by the total number of individuals taken into consideration (with all types of genotypes)

now we know, that the total number of progeny into consideration is 185

BB genotype number is 100

Bb = 75

bb = 10

Therefore

Answer 1 ) The genotype frequency of BB is 100/185 = 0.540

Answer 2) The genpotype frequency of Bb is 75/185 =0.405

Answer 3) The genotype frequency of bb is 10/185 = 0.054

now lets see the indiviidual allele frequency

B = 100*2 (BB genotype ) + 75 (Bb genotype ) = 275

b= 75 (Bb genotype ) + 10*2 (bb genotype ) = 95

Now the total number of genotypes into consideration will be 275+95= 370

Answer 4) Frequency of B allele = 275/370 = 0.743

Answer 5) Frequency of b allele = 95/370 = 0.256

i hope i was able to help you with the answers ! :)